OpenStudy (anonymous):

tan(3x)(tanx-1)=0 [0,2pi)

OpenStudy (loser66):

Where are you stuck?

OpenStudy (anonymous):

I get 0, pi/4, pi/6, 5pi/4, are there other solutions?

OpenStudy (loser66):

pi/3,

OpenStudy (loser66):

4pi/3

OpenStudy (loser66):

7pi/4

OpenStudy (anonymous):

why? then it would be -1

OpenStudy (loser66):

3pi/4

OpenStudy (loser66):

tan (3x) (tanx-1) =0 either tan(3x) =0 or tanx -1 =0, right?

OpenStudy (loser66):

now, tan (3x) =0 iff sin (3x) =0 , ok?

OpenStudy (anonymous):

tanx-1=0

OpenStudy (loser66):

sin ( 3x) =0 when 3x =0 or 3x =pi --> x = pi/2

OpenStudy (loser66):

* x =pi/3

OpenStudy (anonymous):

I get the 3x, but if tanx-1 = 0 then tanx =1 right?

OpenStudy (loser66):

yes,

OpenStudy (anonymous):

so that happens at 45 degrees or 225 degrees

OpenStudy (anonymous):

nevermind...I see it now

OpenStudy (loser66):

not just that, \( tan = \dfrac {sin}{cos}=1 \rightarrow, sin =cos\)

OpenStudy (anonymous):

Thanks

OpenStudy (loser66):

ok

OpenStudy (loser66):

I love helping students who knows what he is doing and willing to fight back if I am wrong.

OpenStudy (anonymous):

:)